# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

### v3

@@ Abstract
Recent advances in large reasoning models highlight Reinforcement Learning with Verifiable Rewards (RLVR) as a promising method for enhancing AI’s capabilities, particularly in solving complex logical tasks. However, it remains unclear whether the current practice of RLVR truly expands a model’s reasoning boundary or mainly amplifies high-reward outputs that the base model already knows for improved precision. This study presents an empirical investigation that provides fresh insights into the potential limits of the common practice of RLVR. We examine how, under current training conditions, RLVR can operate as a support-constrained optimization mechanism that may restrict the discovery of entirely original solutions, remaining constrained by the base model’s initial distribution. We also identify an entropy–reward trade-off: while the current RLVR recipe reliably enhances precision, it may progressively narrow exploration and potentially overlook correct yet underrepresented solutions. Extensive empirical experiments validate that while the current RLVR recipe consistently improves pass@1, the shrinkage of empirical support generally outweighs the expansion of empirical support under larger sampling budgets, failing to recover correct answers that were previously accessible to the base model. Interestingly, we also observe that while RLVR sometimes increases token-level entropy, it results in greater uncertainty at each generation step and declining answer-level entropy. This indicates that these seemingly more uncertain paths ultimately converge onto a smaller set of distinct answers. Taken together, we reveal potential limits of the current RLVR recipe in extending reasoning horizons. Breaking this invisible leash may require future algorithmic innovations such as explicit exploration mechanisms or hybrid strategies that seed probability mass into underrepresented solution regions.
@@ 1 Introduction
The rise of large reasoning models, such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech and others, 2024), marks a breakthrough in AI capabilities, particularly in solving complex logical tasks involving mathematics (Luo et al., 2025c; Zeng et al., 2025) and programming (Luo et al., 2025b; Liu and Zhang, 2025). The key ingredient behind this remarkable progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), where a pretrained base model or one fine-tuned on Chain-of-Thought data is optimized via RL using simple, automatically computed rewards. Prior work has also explored stronger notions of verifiability, such as models that can generate interactive proofs of their own correctness (Amit et al., 2025). Despite empirical success, a fundamental question remains under active debate within the research community: Does the current practice of RLVR expand base models’ reasoning capabilities, or simply reinforce patterns base models already know, sometimes at the expense of exploring alternative correct solutions?
@@
Recent studies have revealed a puzzling pattern that hints at this limitation. While models trained with the common RLVR recipe consistently outperform base models when evaluated with a single attempt (pass@1), base models often perform better given multiple attempts. It is even reported that RLVR benefits from seemingly random or spurious reward signals, raising questions about whether observed improvements genuinely reflect enhanced reasoning (Shao et al., 2025). While pass@k may have limitations as a comprehensive measure of reasoning boundaries, as it primarily captures solution retrieval rather than novel reasoning capacity (Wen et al., 2025), we adopt it here as a practical proxy metric following Chen et al. (2021); Shao and others (2024); Chen et al. (2025b). This metric provides a useful lens for examining how RLVR affects solution accessibility, though future work should explore more nuanced measures of reasoning capability expansion. Besides, prior work examines RLVR only through before/after snapshots, leaving unexplored how RLVR reshapes the model’s effective reasoning support throughout training.
Some studies interpret pass@k as evidence that the current RLVR recipe primarily performs conservative optimization within the base model’s existing capabilities (Yue et al., 2025a; Zhao and others, 2025; Shah and others, 2025; Ma et al., 2025; He et al., 2025). Others argue that this pattern only appears in specialized domains where base models were already well-trained, and that RLVR can substantially expand reasoning in other domains (Liu et al., 2025).
@@
Seeking a definitive answer to this debate remains an open challenge. In the extreme case, it seems unlikely that the current RLVR recipe can unlock advanced reasoning capabilities for any model out of the box, such as GPT-2 (Radford et al., 2019). We are curious if there may exist inherent limitations in the current RLVR practice. This paper provides a systematic empirical investigation into the fundamental capabilities and potential limitations of the current RLVR practice. We introduce the concept of empirical support: the set of correct solutions that a model can realistically discover under finite sampling. Using this framework, we show that:
@@
    The current RLVR recipe primarily preserves rather than expands the base model’s solution coverage. Across diverse reasoning benchmarks, RLVR consistently loses access to more correct solutions than it gains, even while improving single-sample accuracy. We additionally analyze the temporal evolution of support dynamics across RLVR training steps, uncovering how RLVR progressively narrows the accessible solution space over time.
@@
    The precision-diversity trade-off is fundamental, not domain-specific. This pattern appears across mathematics, logical reasoning, factual QA, and code generation, suggesting it reflects inherent properties of current RLVR methods rather than domain-specific quirks.
@@
    Local uncertainty and global diversity can diverge. The current RLVR recipe sometimes increases token-level entropy (appearing more “uncertain” during generation) while simultaneously reducing answer-level entropy (converging to fewer final solutions).
These findings show that current RLVR may face an “invisible leash”. They remain fundamentally constrained by their initialization and cannot discover reasoning patterns that lie outside the base model’s effective reach. To break this invisible leash, RLVR may need to be augmented with explicit exploration or hybrid strategies that seed probability mass into underrepresented regions of the solution space. We hope this work offers novel insights into the strengths and limitations of the current RLVR recipe, guiding future efforts in improving RLVR practice and building LLM systems that can unlock genuinely new reasoning capacity.
2 Preliminaries
@@ 2.1 Formalizing Solution Accessibility
Let $\mathcal{X}$ denote the space of natural language prompts, and $\mathcal{Y}$ denote the space of token sequences (e.g., reasoning traces or completions). For a fixed prompt $x\in\mathcal{X}$, $q(y\mid x)$ is the output distribution of the base model, and $R(x,y)\in\{0,1\}$ is a verifiable reward function indicating whether $y$ is a correct solution. Various RLVR algorithms, including PPO (Schulman et al., 2017), RLOO (Kool et al., 2019), GRPO (Guo et al., 2025), DAPO (Yu et al., 2025), or REINFORCE++ (Hu, 2025), learn a new distribution $\pi_{\theta}(y\mid x)$ to optimize different variants of the following regularized objective: $\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x),x\sim\mathcal{D}}\left[R(x,y)-\beta^{-1}\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\right],$ where $\mathcal{D}$ is the distribution of prompts. An optional log ratio corresponds to a regularized policy update that penalizes divergence from the base model $q$ controlled by a hyperparameter $\beta>0$.
@@
The effective support assumes that $q$ has exact zeros in its support, which, however, rarely holds in practice. Softmax layers yield strictly positive probabilities across all tokens, making the nominal support of $q$ span the entire space $\mathcal{Y}$. This factor, along with sampling noise or temperature scaling, contributes to what we refer to as empirical support diffusion: over time, the model may assign growing probability mass to completions that initially had negligible—but still nonzero—probability under the base model.
@@
2.2 Characterizing How RLVR Changes Solution Access
@@
For a given threshold $\epsilon>0$, We say RLVR achieves empirical support expansion under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(\pi_{\theta})\setminus\mathrm{supp}_{\epsilon}(q)\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that $q(y^{*}\mid x)\leq\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)>\epsilon.$ That is, the RLVR-trained model assigns non-negligible probability mass to correct completions that were effectively negligible under the base model. We say RLVR exhibits empirical support shrinkage under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(q)\setminus\mathrm{supp}_{\epsilon}(\pi_{\theta})\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that $q(y^{*}\mid x)>\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)\leq\epsilon.$ This formalizes the phenomenon where RLVR concentrates probability mass onto a narrower subset of outputs, effectively excluding correct solutions that were previously accessible under the base model.
@@
To quantify RLVR’s impact on solution accessibility, we introduce the following precision and recall-inspired metrics based on these four support categories.
@@
Let $P$, $E$, $S$, and $O$ denote the number of correct completions in preservation, expansion, shrinkage, and out-of-support, respectively. Support Retention Rate (SRR) measures how well RLVR preserves the base model’s accessible correct solutions: $\text{SRR}=\frac{P}{P+S}$ Net Discovery Rate (NDR) measures the fraction of RLVR’s accessible solutions that represent genuine discoveries: $\text{NDR}=\frac{E}{P+E}$ Support Dynamic Score (SDS) provides a balanced measure combining retention and discovery: $\text{SDS}=\frac{2\cdot\text{SRR}\cdot\text{NDR}}{\text{SRR}+\text{NDR}}=\frac{2PE}{P^{2}+2PE+ES}$ Net Support Change Rate (NSCR) captures the net expansion or shrinkage of empirical support: $\text{NSCR}=\frac{E-S}{P+E+S}$
@@ 3 Evidence of Hidden-Support Dynamics
3.1 Experimental Setup
We adopt ProRL-1.5B-v1 (Liu et al., 2025) as our main RLVR method due to its robust long-horizon training framework. Starting from DeepSeek-R1-Distill-Qwen-1.5B as the base model, ProRL’s Nemotron-Research-Reasoning series leverages GRPO enhanced with decoupled clipping, dynamic sampling, KL divergence regularization, and periodic reference resets to sustain exploration and prevent entropy collapse during extended RL training. In addition, we evaluate other RLVR variants at multiple scales (7B–14B parameters), including Skywork (Wei et al., 2023), AceReason-Nemotron (Chen et al., 2025a), and Phi4-Reason (Abdin et al., 2025), alongside a visual LLM (Kangheng-OVR-7B (Wei et al., 2025)).
Table 1: Aggregate support dynamics across diverse models and domains. Each completion is categorized by correctness and support status: Preservation indicates both base and RLVR find the solution; Shrinkage indicates only the base model found it; Expansion indicates only RLVR found it; and Out of Support denotes solutions found by neither. Higher SRR, NDR, and SDS reflect stronger preservation, genuine discovery, and balanced optimization, respectively. NSCR values closer to zero indicate more balanced support change. Kangheng-OVR-7B is included as a vision-language model (VLM). Full detailed statistics for each model are provided in Appx. A.
@@ 3.1 Experimental Setup
OLMo-2-0425-1B Math 0.887 0.110 0.166 -0.072 761 83 104 599
Kangheng-OVR-7B (VLM) Math 1.00 0.00 0.01 -0.00 781 3 4 516
Figure 1: Typical empirical support preservation in Reasoning Gym tasks, like Graph Coloring, Palindrome Generation, and Advanced Geometry.
@@
Figure 3: Examples of empirical support shrinkage on Reasoning Gym tasks such as Leg Counting, Family Relationships, and Power Function.

### v4

@@ Abstract
Recent advances highlight Reinforcement Learning with Verifiable Rewards (RLVR) as a promising method for enhancing LLMs’ capabilities. However, it remains unclear whether the current practice of RLVR truly expands a model’s reasoning boundary or mainly amplifies high-reward outputs that the base model already knows, thereby improving precision. This study presents an empirical investigation that provides fresh insights into the limits of RLVR. We examine how RLVR can operate as a support-constrained optimization mechanism that may restrict the discovery of entirely original solutions, remaining constrained by the base model’s initial distribution. We also identify an entropy-reward trade-off: while RLVR reliably enhances precision, it may progressively narrow exploration and potentially overlook correct yet underrepresented solutions. Extensive empirical experiments validate that while RLVR consistently improves pass@1, the shrinkage of empirical support generally outweighs the expansion of empirical support under larger sampling budgets, failing to recover correct answers that were previously accessible to the base model. Interestingly, while RLVR sometimes increases token-level entropy, it results in greater uncertainty at each generation step and declining answer-level entropy. This indicates that these seemingly more uncertain paths ultimately converge onto a smaller set of distinct answers. Taken together, we reveal potential limits of RLVR in extending reasoning horizons. Breaking this invisible leash requires future innovations that seed probability mass into underrepresented solution regions.
@@ 1 Introduction
Figure 1: Conceptual illustration of empirical support. We define four regions based on whether a correct completion $y^{*}\in\mathcal{C}$ is assigned non-negligible probability mass by the base and RLVR models, $q$ and $\pi_{\theta}$ Support Preservation covers completions with $q(y^{*}|x)>\epsilon$ and $\pi_{\theta}(y^{*}|x)>\epsilon$; Support Shrinkage includes correct completions downweighted by RLVR below $\epsilon$; Support Expansion includes completions that RLVR newly upweights above $\epsilon$ despite negligible base model mass; and Out of Support refers to completions missed by both.
@@
The rise of large reasoning models, such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech and others, 2024), marks a breakthrough in AI capabilities, particularly in solving mathematics (Luo et al., 2025c; Zeng et al., 2025) and programming (Luo et al., 2025b; Liu and Zhang, 2025). The key ingredient behind this remarkable progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), where a pretrained base model is optimized via RL using simple, automatically computed rewards. Prior work has also explored stronger notions of verifiability, such as models that can generate interactive proofs of their own correctness (Amit et al., 2025). Despite empirical success, a fundamental question remains under active debate within the research community: Does the current practice of RLVR expand base models’ reasoning capabilities, or simply reinforce patterns base models already know, sometimes at the expense of exploring alternative correct solutions?
Recent studies have revealed a puzzling pattern that hints at this limitation. While models trained with RLVR consistently outperform base models on a single attempt, base models often perform better with multiple attempts. Shao et al. (2025) also suggests that RLVR can even benefit from noisy or spurious reward signals, further calling into question whether observed gains reflect deeper reasoning improvements or merely sharper exploitation of existing heuristics.
@@
While pass@k may have limitations as a comprehensive measure of reasoning boundaries, as it primarily captures solution retrieval rather than novel reasoning capacity (Wen et al., 2025), we adopt it here as a proxy metric following Chen et al. (2021); Shao and others (2024); Chen et al. (2025b). pass@k provides a useful lens for examining how RLVR affects solution accessibility, though future work should explore more nuanced measures of reasoning capability expansion. Besides, prior work examines RLVR only through before/after snapshots, leaving unexplored how RLVR reshapes the model’s effective reasoning support throughout training. Some studies interpret pass@k as evidence that RLVR primarily performs conservative optimization within the base model’s existing capabilities (Yue et al., 2025a; Zhao and others, 2025; Shah and others, 2025; Ma et al., 2025; He et al., 2025). Others argue that this pattern only appears in specialized domains where base models were already well-trained, and that RLVR can substantially expand reasoning in other domains (Liu et al., 2025a).
@@
Seeking a definitive answer to this debate remains an open challenge. In the extreme case, it seems unlikely that RLVR can unlock advanced reasoning capabilities for any model out of the box, such as GPT-2 (Radford et al., 2019). We are curious if there may exist inherent limitations in the current RLVR practice. This paper provides a systematic empirical investigation into the fundamental capabilities and potential limitations of RLVR. We introduce the concept of empirical support: the set of correct solutions that a model can realistically discover under finite sampling. Using this framework, we show that:
@@
    The current RLVR recipe primarily preserves rather than expands the base model’s solution coverage. Across diverse reasoning benchmarks, RLVR consistently loses access to more correct solutions than it gains, even while improving single-sample accuracy. We also analyze the temporal evolution of support dynamics during RLVR training, revealing how RLVR progressively narrows the accessible solution space.
@@
    The precision-diversity trade-off is fundamental, not domain-specific. This pattern appears across mathematics, logical reasoning, factual QA, and code generation, suggesting it reflects inherent properties of current RLVR rather than domain-specific quirks.
@@
    Local uncertainty and global diversity can diverge. RLVR sometimes increases token-level entropy (appearing more “uncertain” during generation) while simultaneously reducing answer-level entropy (converging to fewer final solutions).
These findings show that RLVR may face an “invisible leash”. They remain fundamentally constrained by their initialization and cannot discover reasoning patterns beyond the base model’s effective reach. To break this invisible leash, RLVR may need to be augmented with explicit exploration or hybrid strategies that seed probability mass into underrepresented regions of the solution space. We hope this work offers novel insights into the strengths and limitations of the current RLVR recipe, guiding future efforts to improve RLVR and build LLM systems that unlock genuinely new reasoning capacity.
2 Related Works
RLVR has emerged as a scalable alternative to RLHF by replacing human preference labels with automated, verifiable signals. Unlike RLHF, which relies on learned reward models for subjective alignment, RLVR optimizes objective, algorithmically computable rewards that directly reflect ground-truth success criteria in domains with clear verifiability. It has achieved strong gains on real-world applications, including math (Guo et al., 2025) and medical multiple-choice QA (Zhang et al., 2025; Lai et al., 2025; Feng et al., 2025), often outperforming SFT with limited training data.
Recent analyses have examined both the benefits and limitations of RLVR. Studies (Wen et al., 2025) argue that standard evaluation metrics, such as Pass@K, can overstate RLVR gains by ignoring the correctness of intermediate reasoning, motivating alternatives (e.g., CoT-Pass@K) that jointly assess reasoning chains and final answers. Under such metrics, RLVR exhibits stronger incentives toward correct reasoning. Complementarily, a recent work by Amit et al. (2025) reports a very interesting theoretical perspective, in which verifier-based training can provide per-instance correctness guarantees by conditioning acceptance on externally verifiable proofs, shifting evaluation from average-case accuracy to worst-case soundness. Concurrent position highlights hidden costs in measurement protocols, data contamination, and evaluation bias, advocating tax-aware evaluation that jointly optimizes accuracy, grounding, and calibrated abstention (Tu et al., 2025).
At the algorithmic level, recent RLVR variants address challenges in reward sparsity, verifier reliability, and exploration. Cai et al. (2025) explicitly models asymmetric noise in automated verifiers to improve robustness under imperfect rewards. Self-verification approaches (Liu et al., 2025b) integrate generation and critique within a single RL loop to produce denser and more informative feedback. Other work explores efficiency-oriented objectives and supervised surrogates to stabilize optimization under sparse verifiable signals (Sun et al., 2025). Finally, emerging theoretical perspectives investigate semantic hidden-state dynamics beyond token-level entropy, proposing structural incentives such as VERL to better decouple exploration and exploitation in reasoning models (Huang et al., 2025).
3 Numerical Metrics
@@ 3.1 Formalizing Solution Accessibility
Let $\mathcal{X}$ denote the space of natural language prompts, and $\mathcal{Y}$ denote the space of token sequences (e.g., reasoning traces or completions). For a fixed prompt $x\in\mathcal{X}$, $q(y\mid x)$ is the output distribution of the base model, and $R(x,y)\in\{0,1\}$ is a verifiable reward function indicating whether $y$ is a correct solution. Various RLVR algorithms, including PPO (Schulman et al., 2017), RLOO (Kool et al., 2019), GRPO (Guo et al., 2025), DAPO (Yu et al., 2025), or REINFORCE++ (Hu, 2025), learn a new distribution $\pi_{\theta}(y\mid x)$ to optimize different variants of the following regularized objective: $\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x),x\sim\mathcal{D}}\left[R(x,y)-\beta^{-1}\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\right],$ where $\mathcal{D}$ is the distribution of prompts. An optional log ratio corresponds to a regularized policy update that penalizes divergence from $q$ controlled by a hyperparameter $\beta>0$.
@@
We define empirical support over answer-level correct completions, where multiple reasoning traces that yield the same verified outcome are treated as a single equivalence class.
@@
The effective support assumes $q=0$, which rarely holds. Softmax layers yield strictly positive probabilities across all tokens, making the nominal support of $q$ span the entire space $\mathcal{Y}$. This factor, along with sampling noise or temperature scaling, contributes to what we refer to as empirical support diffusion: over time, the model may assign growing probability mass to completions that initially had negligible but still nonzero probability under the base model.
@@
3.2 Characterizing Solution Access Change
@@
For a given threshold $\epsilon>0$,
  • We say RLVR achieves empirical support expansion under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(\pi_{\theta})\setminus\mathrm{supp}_{\epsilon}(q)\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that
    $q(y^{*}\mid x)\leq\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)>\epsilon.$
    That is, the RLVR-trained model assigns non-negligible probability mass to correct completions that were effectively negligible under the base model.
  • We say RLVR exhibits empirical support shrinkage under threshold $\epsilon$ if $\mathrm{supp}_{\epsilon}(q)\setminus\mathrm{supp}_{\epsilon}(\pi_{\theta})\neq\emptyset$, i.e. there exists at least one completion $y^{*}\in\mathcal{C}$ such that
    $q(y^{*}\mid x)>\epsilon\quad\text{but}\quad\pi_{\theta}(y^{*}\mid x)\leq\epsilon.$
    This formalizes the phenomenon in which RLVR concentrates probability mass on a narrower subset of outputs, effectively excluding correct solutions that were previously accessible under the base model.
@@
To quantify RLVR’s impact on solution accessibility, we introduce the precision and recall-inspired metrics based on the four support categories.
@@
Let $P$, $E$, $S$, and $O$ denote the number of correct completions in preservation, expansion, shrinkage, and out-of-support, respectively.
  • Support Retention Rate (SRR) measures how well RLVR preserves the base model’s accessible correct solutions:
    $\text{SRR}=\frac{P}{P+S}$
  • Net Discovery Rate (NDR) measures the fraction of RLVR’s accessible solutions that represent genuine discoveries:
    $\text{NDR}=\frac{E}{P+E}$
  • Support Dynamic Score (SDS) provides a balanced measure combining retention and discovery:
    $\text{SDS}=\frac{2\cdot\text{SRR}\cdot\text{NDR}}{\text{SRR}+\text{NDR}}=\frac{2PE}{P^{2}+2PE+ES}$
  • Net Support Change Rate (NSCR) captures the net expansion or shrinkage of empirical support:
    $\text{NSCR}=\frac{E-S}{P+E+S}$
@@ 4 Evidence of Hidden-Support Dynamics
Table 1: Aggregate support dynamics across diverse models and domains. Each completion is categorized by correctness and support status: Preservation indicates both base and RLVR find the solution; Shrinkage indicates only the base model found it; Expansion indicates only RLVR found it; and Out of Support denotes solutions found by neither. Higher SRR, NDR, and SDS reflect stronger preservation, genuine discovery, and balanced optimization, respectively. NSCR values closer to zero indicate more balanced changes in support. Kangheng-OVR-7B is a vision-language model (VLM). Full detailed statistics for each model are in Appx. A.
@@
OLMo-2-0425-1B Math 0.88 0.10 0.18 -0.02 761 83 104 599
Kangheng-OVR-7B Math 1.00 0.00 0.01 -0.00 781 3 4 516
4.1 Experimental Setup
We adopt ProRL-1.5B-v1 (Liu et al., 2025a) as our main RLVR method due to its robust long-horizon training framework. Starting from DeepSeek-R1-Distill-Qwen-1.5B as the base model, ProRL’s Nemotron-Research-Reasoning series leverages GRPO enhanced with decoupled clipping, dynamic sampling, KL divergence regularization, and periodic reference resets to sustain exploration and prevent entropy collapse during extended RL training. In addition, we evaluate other RLVR variants at multiple scales (7B–14B parameters), including Skywork (Wei et al., 2023), AceReason-Nemotron (Chen et al., 2025a), and Phi4-Reason (Abdin et al., 2025), alongside a visual LLM (Kangheng-OVR-7B (Wei et al., 2025)). We additionally include ProRL-1.5B-v2 as a separate checkpoint for comparison (Tab. 1).
Figure 2: Empirical support preservation in Reasoning Gym tasks, like Graph Coloring, Palindrome Generation, and Advanced Geometry.
@@ 4.1 Experimental Setup
Figure 4: Examples of empirical support shrinkage on Reasoning Gym, like Leg Counting and Family Relationships, and Power Function.